OpenStudy (anonymous):

help again on substitutions! Please! 4y-x=5+2y 3x+7y=24

OpenStudy (radar):

x=4y-2y-5=2y-5 substitute for x in 3x+7y=24 3(2y-5)+7y=24 6y-15+7y=24 13y=39 y=3 x=2y-5 x=6-5=1 Correct answer is: y=3 and x=1

OpenStudy (anonymous):

You won't believe where i messed up - I subtracted 15 from 24 instead of adding (6y-15+7y=24) aaaah! Thank you!!!

OpenStudy (radar):

Glad we agree, good luck with your studies.